I already did the first worksheet, i need the second worksheet that depends on the first page i did
Statisties Worksheet 1 Due at the beginning of class on Tuesday, January 29°. This worksheet will count as two online homework grades 1.) Here are the winning margins for the first 42 Super Bowl games (for example the 25 means that the winning team scored 25 more points than the losing team in that Super Bowl): (Already entered in In Class Data Sets) 25, 19,9,16,3,21,7,17,10,4,18, 17,4,12,17,5, 10,29,22, 36 19,32, 4,45, 1.13,35, 17,23,10, 14,7,15,7,27,3,27,3,3,11, 12,3 a.) Form a table showing the class limits, class boundaries, frequency, and relative frequency with 6 classes Class Limits Class Boundaries Frequency Relative Frequency Class width 3 o 3 -IS 1.5. -31|23.5-31.5 | 4 |Olo 049 3.5 007 441 34.5 .5 0:00 b.) Find the mean and standard deviation using StatCrunch. Mean 63S lo bu3
c.) Create a labeled histogram by hand. Use the class boundaries as a scale for the s-asxis Class d.) What is the basic shape: uniform, normal (bell shaped), skewed left, or skewed right? e.) Calculate the coefficient of variation for the Superbowl margins. (Round the mean and standard deviation to at least4 decimal places before calculating the coefficient of variation). f) Find the 5 number summary of the Superbowl margins using StatCrunch. g.) Draw a labeled boxplot of the Superbowl margins.
Answer #1

SolutionC:

12 10 10 4 0) 16 24 32 40 48 Histogram (Frequency Diagram)Solutiond:

positively skewed.

skewed to right.

Solutione:

mean=sum /total=632/42=15.0476

sd=10.5251

coefficient of variation=sd/mean*100

=10.5251/15.04762)*100

=69.95%

Solutionf:

Fivenum summary is

minimum=1

Q1=7

Q2=13.5

Q3=21

maximum=45

Solutiong:

Boxplot S0 45 40 35 30 25 20 15 10 045 36 Plot Area X15.04761905 13.5 5

there is an outlier 45 seen from boxplot